# Initializing Services in Interactive ML Systems for Diverse Users

### v1

@@ Abstract
This paper studies ML systems that interactively learn from users across multiple subpopulations with heterogeneous data distributions. The primary objective is to provide specialized services for different user groups while also predicting user preferences. Once the users select a service based on how well the service anticipated their preference, the services subsequently adapt and refine themselves based on the user data they accumulate, resulting in an iterative, alternating minimization process between users and services (learning dynamics). Employing such tailored approaches has two main challenges: (i) Unknown user preferences: Typically, data on user preferences are unavailable without interaction, and uniform data collection across a large and diverse user base can be prohibitively expensive. (ii) Suboptimal Local Solutions: The total loss (sum of loss functions across all users and all services) landscape is not convex even if the individual losses on a single service are convex, making it likely for the learning dynamics to get stuck in local minima. The final outcome of the aforementioned learning dynamics is thus strongly influenced by the initial set of services offered to users, and is not guaranteed to be close to the globally optimal outcome. In this work, we propose a randomized algorithm to adaptively select very few users to collect preference data from, while simultaneously initializing a set of services. We prove that under mild assumptions on the loss functions, the expected total loss achieved by the algorithm right after initialization is within a factor of the globally optimal total loss with complete user preference data, and this factor scales only logarithmically in the number of services. Our theory is complemented by experiments on real as well as semi-synthetic datasets.
@@ 1 Introduction
In various applications including federated learning (Li et al., 2020), crowd-sourcing (Steinhardt et al., 2016), online recommendation systems (Song et al., 2014), data about user preferences is acquired through interactions with multiple users or sources. The primary objective is to leverage this data to deliver services that cater to the diverse requirements of users. Typically, these users exhibit heterogeneity in their preferences, and recent approaches demonstrate the benefits of deploying and updating multiple, more specialized models versus using one large shared model as a one-size fits all approach; e.g., for clustered federated learning Mansour et al. (2020); Sattler et al. (2020); Ghosh et al. (2020) (Dwork et al., 2018), in the context of fair classifiers (Ustun et al., 2019), and for meta learning (Kong et al., 2020). And even when a single large model is used, it is often followed by a fine-tuning step for specialization to a group of users or tasks by retraining part of the model (Chua et al., 2021; Sun et al., 2021).
In this paper we consider ML systems that interactively learn from users across multiple subpopulations with heterogeneous data distributions. The aim is to provide a range of services, each service specialized for a group of users. These services typically work by trying to predict user preferences, and the likelihood of any user picking a service from the available candidates depends on the services’ performance for them. For example, Meta [^1] https://nickclegg.medium.com/you-and-the-algorithm-it-takes-two-to-tango-7722b19aa1c2 uses ML to predict user preferences and make personalized recommendations for its services like Facebook, Instagram, etc. Users decide whether to join recommended pages, leave likes or other forms of engagement such as scrolling activity, hours online indicating their level of satisfaction with the services. Meta would in turn plan which users each of its platforms is tailored to, and what personalized content is recommended to users—these decisions by Meta could altogether be modeled as choosing the parameters of its services.
These services are usually trained through interaction with users. Once a set of services is launched users select the service that best predicts their preferences (generally choosing the service that minimizes the particular user’s performance metric). These services are then subsequently refined based on data of current users. This sequence of (alternating) decisions by users and by services (learning dynamics) continues as more data is gathered. Broadly speaking, such a learning system could include a set of independent learners (e.g., different service providers) having an incentive to coordinate for the purpose of initializing their services, or a single learner with several deployed models (e.g., a company that owns different platforms, or a multi-model server in certain federated learning setups). This setup has been studied by Ginart et al. (2021), who discuss the effect of competition between services on user welfare, and Dean et al. (2022), who discuss the the stability of equilibria of learning dynamics in such systems.
Our goal is to design a set of services to minimize the total loss, i.e., the sum of all losses across users. However, we face the following challenges:
@@
  • Unknown User Preferences: In practice, the presence of a service often precedes data collection. Specifically, in contexts like online recommendations, it is usually not feasible to gather user preference data without first deploying the service and observing user interactions. This means that data collection is inherently conditional on the existence of a service, challenging the conventional “data-first, model-second” paradigm.
  • Suboptimal Local Solutions: The total loss (sum of losses across all users and all services) landscape in these applications is non-convex (jointly in both sets of variables for users and services) even if the individual user losses on a single service are convex, making it likely for the learning dynamics to get stuck in local minima where the total loss can be significantly worse than the globally minimum loss. Thus, the outcomes of learning dynamics are heavily affected by initialization.
Contributions.
The following summarizes the contributions of this paper.
@@
  • In this work, we address these challenges by designing a computationally and statistically efficient algorithm for initialization of services prior to learning dynamics. The algorithm works by adaptively selecting very few users to collect data from (via zeroth order or bandit access to their loss functions) in order to initialize the set of services.
@@
  • We establish an approximation ratio for our algorithm: the expected total loss achieved by the algorithm right after initialization is within a factor of the globally optimal total loss in the presence of complete user preference data, and this factor scales only logarithmically in the number of services (cf. Section 3).
@@
  • When users belong to a set of unknown demographic groups, it is desirable that the provided services do not result in unfavorable outcomes towards certain demographics (say based on gender or racial groups). A fair objective would be to minimize the maximum average loss of users across different groups. We provide an approximation ratio for our algorithm on the fair objective as well, that scales only logarithmically in the number of services (cf. Section 3).
@@
  • In the context of linear prediction models, we study generalization properties of the initialized services to users the services have never interacted with before (cf. Section 4).
@@
  • We empirically demonstrate the strengths of our initialization scheme through experiments on a prediction task based on 2021 US Census data, and online movie recommendation task based on Movielens10M dataset (cf. Section 5).
Notation and Terminology.
@@
For a symmetric matrix $\mathbf{A}$ and any vector $x\in\mathbb{R}^{d}$, we denote its Mahalanobis norm by $\|x\|_{\mathbf{A}}=\sqrt{x^{\top}\mathbf{A}x}$. The generalized eigenvalues for a pair of symmetric matrices $\mathbf{A}$ and $\mathbf{B}$ are denoted by $\lambda(\mathbf{A},\mathbf{B})$, defined as the solutions of $\lambda$ for the generalized eigenvalue problems $\mathbf{A}v=\lambda\mathbf{B}v$ (Ghojogh et al., 2019; Förstner and Moonen, 2003). Specifically we use $\lambda_{\min}(\mathbf{A},\mathbf{B})$ to denote minimum generalized eigenvalue for the matrix pair $\mathbf{A},\mathbf{B}$.
The loss for a user $i$ given service $\theta\in\mathbb{R}^{d}$ is denoted by $\mathcal{L}_{i}(\theta,\phi_{i})$ where $\phi_{i}$ parameterizes the user’s preference. For a set of users $\mathcal{A}$ (e.g., $\mathcal{A}=[n]$ denotes a set of $n$ users) and a set of services $\Theta=\{\theta_{1},\ldots,\theta_{k}\}\subset\mathbb{R}^{d}$, the total loss is defined as $\mathcal{L}(\Theta,\mathcal{A})=\sum_{i\in\mathcal{A}}\min_{j\in[k]}\mathcal{L}_{i}(\theta_{j},\phi_{i})$.
@@ 2 Problem Setup
We consider the setting where a provider wants to design $k$ services for $n$ users with unknown diverse preferences, such that each service specializes to a subgroup of users. Each service typically uses a model parameterized by a vector $\theta\in\mathbb{R}^{d}$, which predicts users’ preferences. A user is more likely to use a service that yields higher utility by predicting their preference more accurately (corresponding to lower prediction loss).
Let the loss incurred by some user $i$ upon choosing some service parameterized by $\theta\in\mathbb{R}^{d}$ be denoted by $\mathcal{L}_{i}(\theta,\phi_{i})$. Here $\phi_{i}\in\mathbb{R}^{d}$ represents this user’s unknown preference parameter and $\mathcal{L}_{i}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ is an unknown (nonnegative) function governing this user’s loss (which the provider can only evaluate at a given point; i.e., has zeroth order or bandit access to this function only, see Section 3 for more details).
For example, in the case of movie recommendations, $\phi_{i}$ could be the user’s interests across different genres of movies, $\theta$ a set of recommended movies, and $\mathcal{L}_{i}(\theta,\phi_{i})$ measures user dissatisfaction, say some function of hours spent online for this user, or a rating the user left upon using this recommendation. Users can have diverse preferences and loss geometries. The example in Figure 1 emphasizes that users not only have distinct preferences, but also experience varying degrees of dissatisfaction or satisfaction when presented with recommendations not exactly their preference.
Note that information about $\mathcal{L}_{i}$ and $\phi_{i}$ is obtained only upon interacting with the users, which we will discuss in Section 3.
Figure 1: Imagine two users who both have a strong preference for action movies (and both are content if they receive recommendations for this genre). However, User 1 also enjoys romantic comedies, but dislikes documentaries, while User 2 has the opposite preference. If User 1 is recommended a documentary, they will be dissatisfied, whereas User 2 will still remain content. Consequently, despite their shared fondness for action movies, their levels of dissatisfaction or satisfaction diverge when they encounter content from a different genre.
We make the following assumptions about the functional form of $\mathcal{L}_{i}$, and state several examples of function classes satisfying these properties.
Assumption 1 (Unique Minimizer).
@@
This assumption implies that unless all users have identical preferences, there doesn’t exist a single service parameter $\theta$ that simultaneously minimizes every user’s loss. Thus providing multiple services (multiple $\theta$’s) where the users choose the one best for them is strictly better than one service for all users.
Assumption 2 (Approximate Triangle Inequalities).
@@
Here $c_{ij}$ (equal to $c_{ji}$) captures the alignment between the preferences and loss geometries of two users. Lower values of $c_{ij}$ indicate less similarity. Item $(i)$ implies that the loss for user $i$ on any service $\theta\in\mathbb{R}^{d}$ is no worse than (up to a constant factor) the sum of (a) loss of another user $j$ on using the same service, and (b) the loss of user $j$ if they were to use user $i$’s preference. The latter term (b) can be seen as measuring the similarity between the users’ preferences.
For condition $(ii)$ to hold for all $\theta\in\mathbb{R}^{d}$, it must also hold for the service that minimizes the sum of losses of both users using the same service. Suppose that users $i$ and $j$ were to exchange preferences. Then their loss would be no worse than (up to a constant factor) their minimum total loss, i.e.,
$\mathcal{L}_{i}(\phi_{j},\phi_{i})\leq\frac{1}{c_{ij}}\min_{\theta\in\mathbb{R}^{d}}\left(\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{i}(\theta,\phi_{i})\right).$
Some examples of functions (with corresponding constants) satisfying these assumptions include the following (see Appendix A for additional examples and derivations):
@@
  • The Mahalanobis distance: $\mathcal{L}_{i}(\theta,\phi_{i})=\|\theta-\phi_{i}\|_{\Sigma_{i}}$. Different users can have different $\Sigma_{i}$ capturing their diverse loss variation (cf. Figure 1), as long as $\Sigma_{i}$ is full rank. Here $c_{ij}=\min\{\lambda_{\rm min}(\Sigma_{i},\Sigma_{j}),\lambda_{\rm min}(\Sigma_{j},\Sigma_{i})\}$.
  • Any distance metric: This naturally follows from triangle inequality, hence $c_{ij}=1$. (Note that the examples above are also all distances, but when focusing on their special form we obtain better estimates for the $c_{ij}$.)
Suppose the users have access to $k$ services parameterized by $\Theta=\{\theta_{1},\ldots,\theta_{k}\}\subset\mathbb{R}^{d}$. Then, each user $i$ selects service $\theta_{l}$ that minimizes their loss, i.e. $\mathcal{L}_{i}(\Theta,\phi_{i})=\min_{l\in[k]}\mathcal{L}_{i}(\theta_{l},\phi_{i})$.
Our objective is to design $\Theta$, such that the sum of losses across users and services is minimized. We define the objective as follows:
$\mathcal{L}(\Theta,[n])=\sum_{i\in[n]}\min_{j\in[k]}\mathcal{L}_{i}(\theta_{j},\phi_{i}).$ (1)
@@
Since preferences $\{\phi_{i}\}_{i\in[n]}$ and loss functions $\{\mathcal{L}_{i}\}_{i\in[n]}$ are unknown, and the provider only has zeroth order or bandit feedback access, estimating the objective function $\mathcal{L}(\Theta,[n])$ usually needs a lot of data collected uniformly across the users. This large amount of data is needed before services can be deployed, yet as stated earlier, we are in the situation where we have no data until we deploy services and observe user interactions. Our limited access to user information (via limited queries) make our setting challenging.
@@
In addition to the first challenge, the loss function is non-convex and iterative minimization approaches from a random initialization are susceptible to getting stuck in arbitrarily poor local optima. Alternatively, computing the optimal clustering first and then finding the best service for each cluster would be NP-Hard.
Despite these challenges, in Section 3, we propose an algorithm that is both statistically and computationally efficient, and admits an approximation ratio with respect to the globally optimal value, i.e., $\mathcal{L}(\Theta_{\rm OPT},[n])$.

### v2

@@ Abstract
This paper investigates ML systems serving a group of users, with multiple models/services, each aimed at specializing to a sub-group of users. We consider settings where upon deploying a set of services, users choose the one minimizing their personal losses and the learner iteratively learns by interacting with diverse users. Prior research shows that the outcomes of learning dynamics, which comprise both the services’ adjustments and users’ service selections, hinge significantly on the initialization. However, finding good initializations faces two main challenges: (i) Bandit feedback: Typically, data on user preferences are not available before deploying services and observing user behavior; (ii) Suboptimal local solutions: The total loss landscape (i.e., the sum of loss functions across all users and services) is not convex and gradient-based algorithms can get stuck in poor local minima.
We address these challenges with a randomized algorithm to adaptively select a minimal set of users for data collection in order to initialize a set of services. Under mild assumptions on the loss functions, we prove that our initialization leads to a total loss within a factor of the globally optimal total loss with complete user preference data, and this factor scales logarithmically in the number of services. This result is a generalization of the well-known $k$-means++ guarantee to a broad problem class, which is also of independent interest. The theory is complemented by experiments on real as well as semi-synthetic datasets.
@@ 1 Introduction
We consider a setting where a provider wants to design $k$ services for $n$ users with diverse preferences. Each service uses a model parameterized by a vector $\theta\in\mathbb{R}^{d}$ to predict users’ preferences, and users pick a service that yields the smallest loss for them. The loss incurred by user $i$ when choosing a service parameterized by $\theta$ is denoted by $\mathcal{L}_{i}(\theta,\phi_{i})$, where $\phi_{i}\in\mathbb{R}^{d}$ parameterizes the user’s preference. We want to design $k$ services by minimizing the sum of all losses; i.e., an optimization problem of the form:
$\operatorname*{minimize}_{\theta_{1},\ldots,\theta_{k}\in\mathbb{R}^{d}}\;\;$ $\sum_{i=1}^{n}\min\{\mathcal{L}_{i}(\theta_{1},\phi_{i}),\ldots\mathcal{L}_{i}(\theta_{k},\phi_{i})\}.$ (1)
This problem formulation is broad and includes the classical $k$-means clustering problem [28] (where $\mathcal{L}_{i}$ are Euclidean distances and the inner ‘min’ selects the closest among $k$ centroids), mixed-linear regression [42], generalized principal component analysis (GPCA) or subspace clustering [1], in addition to our new, motivating problem of designing $k$ services for $n$ users. Even if the losses $\mathcal{L}_{i}$ are convex in $\theta$, this objective is generally not convex (even in the special case of the $k$-means problem).
Our goal is to find a local minimum of this optimization problem with an approximation ratio (a worst-case guarantee on the achieved total loss with respect to the global optimum) under suitable yet broad assumptions on $\mathcal{L}_{i}$. Further, an important limitation in many practical settings is that the provider/designer has only bandit feedback (zeroth-order oracle access) to the loss functions (i.e., the designer doesn’t know the function $\mathcal{L}_{i}(\cdot,\phi_{i})$, but can only evaluate its value for some $\theta$ corresponding to the service chosen by the user among the ones deployed), which further complicates solution methods, compared to the classical cases of clustering and facility location problems [10] which typically assume full information.
In this paper, we seek a novel and effective initialization scheme that vastly extends the celebrated $k$-means++ algorithm and its analysis [5]. This scheme should retain the simplicity and ease of implementation of the original algorithm, yet be able to (1) handle general loss families (assumptions on $\mathcal{L}_{i}$ are discussed in Section 2), (2) provide a tight, instance-dependent approximation ratio (details in Section 3), and (3) handle realistic information limitations such as access only to (noisy) bandit feedback, in a sample-efficient manner. Next, we describe in more detail the important use case of initialization of services for diverse users in multi-service ML systems (yet as noted above, our main result has other applications as well, and can be of independent interest).
Motivation. In a variety of contexts such as federated learning [27], crowd-sourcing [36] and online recommendation systems [35], data about user preferences is acquired through iterative interactions. This data is then used to improve the model and serve the individual needs of users. Given that users’ preferences are typically heterogeneous, recent works demonstrate that using multiple specialized models can be more effective than the one-size-fits-all approach of employing a single large shared model, e.g., for clustered federated learning [30, 34, 17], meta learning [25, 7], fine-tuning for specific groups of users or tasks [9, 37], and in the context of fair classifiers [39]. Here we tackle the crucial yet under-explored phase of initializing services in ML systems that learn interactively from diverse users. The initialization process is crucial as it sets the stage for how effectively these systems can adapt and specialize with future user interactions. Once initialized, the services interact with users, who, in turn, choose among services based on their loss. These “learning dynamics” typically lead to the specialization of services to groups of users [18, 12], and [12] shows experimentally that the overall social welfare achieved by the services depends on the initialization of the learning dynamics. We note that in the context of problem (1), the learning dynamics in [12] can be seen as updates of an alternating minimization algorithm, iteratively updating users’ choices (the inner minimization) and services’ parameters (update to each $\theta_{j}$). Our goal is to initialize a set of services to minimize the sum of losses for all users (or equivalently, maximize total welfare), tackling the following challenges:
@@
  • Bandit loss feedback: In practice, offering a service often precedes data collection. Specifically, in contexts like online recommendations, it is usually not feasible to gather user preference data (knowledge about $\phi_{i}$ and evaluations of $\mathcal{L}_{i}(\cdot,\phi_{i})$ at various different parameter values in problem (1)) without first deploying the services parameterized by $\{\theta_{1},\ldots,\theta_{k}\}$ and observing user interactions. This means that data collection is inherently conditional on the existence of services, challenging the conventional “data-first, model-second” paradigm.
  • Suboptimal local solutions: Since users select the service with the lowest loss ($\min\{\mathcal{L}_{i}(\theta_{1},\phi_{i}),\ldots\mathcal{L}_{i}(\theta_{k},\phi_{i})\}$ in problem (1)), minimizing over the parameters $\{\theta_{1},\ldots,\theta_{k}\}$ leads to a nonconvex problem in general, as mentioned earlier. Gradient-based learning dynamics can get stuck in local minima where the total loss can be significantly worse than the globally minimum loss. Thus, the outcomes of learning dynamics are heavily affected by initialization of service parameters.
Contributions. The following summarizes the contributions of this paper.
@@
  • We design a computationally and statistically efficient algorithm for initialization of services prior to learning dynamics. The algorithm works by adaptively selecting a small number of users to collect data from (via queries of their loss function) in order to initialize the set of services.
@@
  • We establish an approximation ratio for the designed algorithm: the expected total loss achieved by the algorithm right after initialization is within a factor of the globally optimal total loss in the presence of complete user preference data, and this factor scales logarithmically in the number of services. Furthermore this bound is tight, and recovers the known k-means++ approximation ratio as a special case (cf. Section 3).
@@
  • When users belong to a set of demographic groups, it is desirable that the services do not result in unfavorable outcomes towards certain demographics (e.g., based on gender or racial groups). One fair objective is to minimize the maximum average loss of users across different groups. We provide an approximation ratio for this fair objective that scales logarithmically in the number of services (cf. Section 3).
@@
  • In the context of linear prediction models, we study the problem of generalizing to users the provider has not interacted with before (cf. Section 4).
@@
  • We empirically demonstrate the strengths of our initialization scheme via experiments on a prediction task using 2021 US Census data, and online movie recommendation task using the Movielens10M dataset (cf. Section 5).
1.1 Related Work
Multiple Model Specialization. In distributed learning, where data sources are users’ personal devices, utilizing multiple specialized models, where users are grouped into clusters representing interests, can yield improved predictions and outcomes. For instance, in recommendation systems these clusters could represent users interested in different movie genres, or different combination of features (see Appendix B for a concrete example on Netflix recommendation clusters). This approach has been adopted recently in clustered federated learning [34, 30, 17] and online interactive learning [32], facility location problems [6], where users choose models/services and for which they provide updates.
Clustering. Multiple model specialization leads to clustering the users into groups and centering a specialized model on each group. We provide a brief review of the $k$-means clustering problem and establish the connection to specialization. The $k$-means clustering problem is one of the most commonly encountered unsupervised learning problems. Given a set of known $n$ points in Euclidean space, the goal is to partition them into $k$ clusters (each characterized by a center), such that the sum of square of distances to their closest center is minimized. Dasgupta [11] and Aloise et al. [3] showed that the $k$-means problem is NP-Hard. The most popular heuristic for $k$-means is Lloyd’s algorithm [28], which proceeds by randomly initializing $k$ centers and then uses iterative updates to find a locally optimal $k$-means clustering, which can be arbitrarily bad compared to the globally optimal clustering. The performance of the $k$-means algorithm relies crucially on the initialization. Arthur and Vassilvitskii [5] and Ostrovsky et al. [33] proposed an elegant polynomial time algorithm for initializing centers, known as $k$-means++. Arthur and Vassilvitskii [5] proved that the expected cost of the initial clustering obtained by $k$-means++ is at most $8(2+\log k)$ times the cost of optimal $k$-means clustering. Our work generalizes the analysis of Arthur and Vassilvitskii [5] to the setting where user’s preferences are represented as unknown points and the loss functions are unknown with only bandit access, not necessarily identical, and general as long they satisfy Assumptions 2.1 and 2.2, with important examples given in Appendix C.
For a detailed discussion on more related works please see Appendix A.
@@ 1.1 Related Work
Notation and Terminology. For a symmetric matrix $\mathbf{A}$ and any vector $x\in\mathbb{R}^{d}$, we denote its Mahalanobis norm by $\|x\|_{\mathbf{A}}=\sqrt{x^{\top}\mathbf{A}x}$. The generalized eigenvalues for a pair of symmetric matrices $\mathbf{A}$ and $\mathbf{B}$ are denoted by $\lambda(\mathbf{A},\mathbf{B})$, defined as the solutions of $\lambda$ for the generalized eigenvalue problems $\mathbf{A}v=\lambda\mathbf{B}v$ [16, 14]. Specifically we use $\lambda_{\min}(\mathbf{A},\mathbf{B})$ to denote minimum generalized eigenvalue for the matrix pair $\mathbf{A},\mathbf{B}$. The loss for a user $i$ given service $\theta\in\mathbb{R}^{d}$ is denoted by $\mathcal{L}_{i}(\theta,\phi_{i})$ where $\phi_{i}$ parameterizes the user’s preference. For a set of users $\mathcal{A}$ (e.g., $\mathcal{A}=[n]$ denotes a set of $n$ users) and a set of services $\Theta=\{\theta_{1},\ldots,\theta_{k}\}\subset\mathbb{R}^{d}$, the total loss is defined as $\mathcal{L}(\Theta,\mathcal{A})=\sum_{i\in\mathcal{A}}\min_{j\in[k]}\mathcal{L}_{i}(\theta_{j},\phi_{i})$.
@@ 2 Problem Setup
We make the following assumptions about the functional form of $\mathcal{L}_{i}$, and state several examples of function classes satisfying these properties. Note that the designer/provider doesn’t need to know the functional form of $\mathcal{L}_{i}$, and knowledge about $\mathcal{L}_{i}$ is obtained through bandit feedback via observing the scalar values $\mathcal{L}_{i}(\theta,\phi_{i})$ for different $\theta$, where $\theta$ parameterizes the services.
Assumption 2.1 (Unique Minimizer).
@@
This assumption implies that unless all users have identical preference parameters, there doesn’t exist a single service parameter $\theta$ that simultaneously minimizes every user’s loss. Thus providing multiple services (multiple $\theta$’s) where the users choose the one best for them is strictly better than one service for all users.
Assumption 2.2 (Approximate Triangle Inequalities).
@@
Here $c_{ij}$ (equal to $c_{ji}$) captures the alignment between the preference parameters and loss geometries of two users. Lower values of $c_{ij}$ indicate less similarity. Item $(i)$ implies that the loss for user $i$ on any service $\theta\in\mathbb{R}^{d}$ is no worse than (up to a constant factor) the sum of (a) loss of another user $j$ on using the same service, and (b) the loss of user $j$ if they were to use user $i$’s preference parameter. The latter term (b) can be seen as measuring the similarity between the users’ preferences.
For condition $(ii)$ to hold for all $\theta\in\mathbb{R}^{d}$, it must also hold for the service that minimizes the sum of losses of both users using the same service. Suppose that users $i$ and $j$ were to exchange preference parameters. Then their loss would be no worse than (up to a constant factor) their minimum total loss, i.e.,
$c_{ij}\mathcal{L}_{i}(\phi_{j},\phi_{i})\leq\min_{\theta\in\mathbb{R}^{d}}\left(\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{i}(\theta,\phi_{i})\right).$
Some examples of loss functions and the corresponding constants that satisfy these assumptions include the following (see Appendix C for additional examples and derivations):
@@
  • The Mahalanobis distance: $\mathcal{L}_{i}(\theta,\phi_{i})=\|\theta-\phi_{i}\|_{\Sigma_{i}}$. Different users can have different $\Sigma_{i}$ capturing their diverse loss variation, as long as $\Sigma_{i}$ is full rank. Here $c_{ij}=\min\{\lambda_{\rm min}(\Sigma_{i},\Sigma_{j}),\lambda_{\rm min}(\Sigma_{j},\Sigma_{i})\}$.
  • Any distance metric: This naturally follows from triangle inequality, hence $c_{ij}=1$.
  • Any arbitrary function $\mathcal{L}_{i}(\theta,\phi_{i})$ that is $L_{i}$-Lipschitz and $\mu_{i}$-strongly convex in $\theta$ with $c_{ij}=\min(\mu_{i},\mu_{j})/\max(L_{i},L_{j})$.
Objective. Suppose the users have access to $k$ services parameterized by $\Theta=\{\theta_{1},\ldots,\theta_{k}\}\subset\mathbb{R}^{d}$. Then, each user $i$ selects service $\theta_{l}$ that minimizes their loss, i.e. $\mathcal{L}_{i}(\Theta,\phi_{i})=\min_{l\in[k]}\mathcal{L}_{i}(\theta_{l},\phi_{i})$.
As discussed earlier, our goal is to design $\Theta$, such that the sum of losses across users and services is minimized. We define the objective as follows:
@@
Since preferences $\{\phi_{i}\}_{i\in[n]}$ and loss functions $\{\mathcal{L}_{i}\}_{i\in[n]}$ are unknown and the provider only has zeroth order or bandit feedback access, estimating the objective function $\mathcal{L}(\Theta,[n])$ usually needs a lot of data collected uniformly across the users. This large amount of data is needed before services can be deployed, yet as stated earlier, we are in the situation where we have no data until we deploy services and observe user interactions. Our limited access to user information (via limited queries) makes our setting challenging.
@@
The loss function is non-convex and iterative minimization approaches from a random initialization are susceptible to getting stuck in arbitrarily poor local optima. This means computing the optimal clustering first and then finding the best service for each cluster is NP-Hard.
Challenge 3.
We do not assume any data separability conditions, for example user preference parameters are drawn from $k$ well separated distributions. Thus we are unable to exploit underlying structure to reduce sample complexity.
Despite the challenges, in Section 3, we propose an algorithm that is both statistically and computationally efficient, and admits an approximation ratio with respect to the globally optimal value, i.e., $\mathcal{L}(\Theta_{\rm OPT},[n])$.